OpenStudy (anonymous):

Find the product: (x + 5)(x − 5) x2 + 25 x2 − 25 x2 − 10 x2 − 5

OpenStudy (anonymous):

x2 − 25 (a+b)(a-b)=a²-b²

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow (a + b)(a - b) = a^2 - b^2 \Longleftrightarrow (x + 5)(x - 5) = x^2 - 5^2 }\) So x^2 - 25

OpenStudy (anonymous):

dude seriously. you keep asking similar questions. >.< Just multiply the answers. :/